# The Coverage Principle: A Framework for Understanding Compositional Generalization

### v1

@@ F.1 Measurement protocol for $N_{\mathrm{req}}$
To empirically determine the minimum dataset size required for reliable compositional generalization ($N_{\mathrm{req}}$), we develop a measurement protocol that accounts for practical computational constraints while ensuring robustness. For each token set size $|\mathcal{X}|$ and task structure, we test multiple dataset sizes until we identify the threshold point where the model successfully generalizes to the ID test set.
@@
    Measurement precision: For each identified $N_{\mathrm{req}}$, we verified that 75% of this dataset size consistently failed to meet our generalization criterion. This establishes that our measurement error is at most $-\log(0.75)=0.125$ in log scale, providing confidence in the derived power-law exponents.
F.2 Measured power-law scaling constants across task structures and model sizes
Using our measurement protocol, we measure the required dataset size $N_{\mathrm{req}}$ across three different compositional structures (2-Hop, Parallel-2-Hop, and 3-Hop) and three model scales (68M, 96M, and 1.5B parameters). For each task structure, we vary the token set size $|\mathcal{X}|$ from 50 to 200, allowing us to observe the scaling relationship.
Table 2 presents the power-law exponents obtained by linear fitting $\log(|\mathcal{X}|)$ vs. $\log(N_{\mathrm{req}})$ plots, all with $R^{2}>0.99$. The consistency of exponents across model sizes suggests that the observed power-law scaling relates to properties of the compositional tasks themselves, rather than model capacity. This observation aligns with our theoretical derivation in Section 5.1, which predicts that the required dataset size scales at least quadratically with token set size.
Table 2: Power law exponents for different tasks and GPT-2 sizes, obtained by linear fitting $\log(|\mathcal{X}|)$ vs. $\log(N_{\mathrm{req}})$ plots. $R^{2}>0.99$ for all linear fitting.
@@ F.2 Measured power-law scaling constants across task structures and model sizes
F.3 Robustness to hyperparameter variations
Figure 15: Robustness of power-law scaling relationship to hyperparameter variations in the 2-Hop task with $|\mathcal{X}|=50$. Each line shows the training and test accuracy curves for a different configuration: (1) baseline, (2) reduced learning rate (4e-4, half of baseline), (3) reduced weight decay (0.01, one-tenth of baseline), and (4) changed generalization criteria (test accuracy > 0.95 within 10 epochs after training accuracy > 0.95). $R^{2}>0.99$ for all linear fitting.
To verify that our observed power-law scaling relationship is not an artifact of specific hyperparameter choices, we conduct ablation studies with modified training configurations. Figure 15 demonstrates that for the 2-Hop task with $|\mathcal{X}|=50$, the following changes did not significantly affect the measured $N_{\mathrm{req}}$ or the derived power-law exponent:
@@ F.3 Robustness to hyperparameter variations
Appendix G Detailed analysis for Non-tree task
This section provides additional analyses that support our findings in Sec. 5.4 regarding the challenges of path ambiguity in the Non-tree task.
G.1 Coverage analysis
Figure 16: Coverage analysis for Non-tree task with $|\mathcal{X}|=50$. The graph shows the percentage of ID test data covered at different $k$ values across various dataset sizes ($N$). Compared to the 2-Hop task (Fig. 3, left), Non-tree has significantly lower coverage at equivalent dataset sizes, indicating that path ambiguity impedes the formation of functional equivalence relationships.
Fig. 16 demonstrates that with equivalent training dataset sizes, a smaller percentage of ID test examples fall inside $k$-coverage for the Non-tree task compared to the 2-Hop task shown in Fig. 3 (Left). This aligns with our theoretical analysis in Sec. 5.4, which predicts that path ambiguity limits the establishment of functional equivalence relationships between input subsequences, as the model cannot generalize across different $x_{2}$ values in the Non-tree structure even when they produce the same intermediate state $b=f_{1}(x_{1},x_{2})$.
G.2 Effect of model scaling
Figure 17: ID test accuracy comparison between GPT-2 (96M parameters) and GPT-2-XL (1.5B parameters) on the Non-tree task with $|\mathcal{X}|=50$, measured 100 epochs after training accuracy exceeds 0.99. Despite the 15x increase in parameter count, the accuracy does not increase.
Fig. 17 shows that scaling up the model size to GPT-2-XL (1.5B parameters) does not significantly improve generalization performance on the Non-tree task, even when measured 100 epochs after reaching training accuracy > 0.99. This suggests that the challenges posed by path ambiguity cannot be overcome simply by increasing model capacity, supporting our claim that the limitation is structural rather than related to model capacity.
G.3 Representation analysis in successful generalization
For a model that eventually achieved near-perfect ID accuracy (0.96) after extended training (36k epochs, $|\mathcal{X}|=50$, $N=50\text{k}$), we conduct causal tracing analysis to understand how it achieves generalization despite path ambiguity.
@@ G.3 Representation analysis in successful generalization
The causal tracing results in Fig. 18 reveal how the model achieves generalization in the presence of path ambiguity. Across all perturbation strategies, the model’s predictions show strong causal dependence on representations at both the $x_{1}$ and $x_{2}$ positions, indicating reliance on direct access to both input tokens rather than an abstracted intermediate computation. This pattern contrasts sharply with the 2-Hop task, where causal effects concentrate primarily at positions corresponding to clustered functional equivalence representations.
@@
Appendix H Detailed discussion on the taxonomy of generalization
In Section 7, we introduce a taxonomy of generalization that extends beyond the coverage principle. We propose this taxonomy as an initial conceptual framework to disentangle different mechanisms of generalization that are often conflated in the literature. This categorization offers hypotheses for why neural networks succeed or fail at compositional tasks, connecting empirical observations to theoretical principles. We emphasize that the taxonomy is preliminary and that Sec. H.5 outlines its current limits, open questions, and future directions.
H.1 Type-I: Structure-based generalization
Type-I generalization occurs when a model succeeds by identifying and utilizing functionally equivalent components based on how primitive functions are composed. This form of generalization is precisely what we have formalized and studied via the coverage principle. Crucially, this kind of generalization remains strictly bounded by coverage as defined in Section 3.
In Type-I generalization, a model learns to recognize when different input fragments yield identical results in shared contexts, allowing it to generalize to new combinations of these fragments. However, without exposure to sufficient functional equivalence evidence during training, the model cannot reliably generalize to novel compositions.
Type-I generalization effectively describes what a pattern matching learner can achieve through the functional equivalence observed by input-output relationships, without requiring abstract variable binding mechanisms. This limitation aligns with the classic argument by Fodor and Pylyshyn [9] and Marcus [37] that connectionist networks lacking explicit variable binding capabilities [^13] We expand on the variable binding concept defined in Sec. 2 beyond flexibly associating variables with values, and we distinguish between self-replicative expansion (e.g., 2-hop $\to$ 3-hop via Type-III) and complex compositional binding (e.g., substituting noun phrases for nouns), where the latter may require explicit variable binding mechanisms beyond parameter sharing. cannot exhibit fully systematic generalization. We introduce this perspective as a conceptual framing of the coverage principle: it describes the ceiling of pattern-based generalization in neural networks that rely on pattern matching rather than symbolic variable binding.
H.2 Type-II: Function property-based generalization
Type-II generalization occurs when a model exploits specific properties intrinsic to individual primitive functions, enabling generalization beyond patterns observed during training. Unlike Type-I generalization which requires explicit evidence of functional equivalence, Type-II generalization leverages invariant properties that hold for all inputs of the primitive, making the function’s output fully determined by those invariances even in unobserved contexts.
We identify several sub-families of property-based generalization:
  • Algebraic invariances: commutativity, associativity, and similar properties
  • Group-theoretic structure: as in Fourier basis representations for modular arithmetic [63]
  • Input irrelevance: attention sparsity that ignores distractors [81]
The reversal curse [16] exemplifies the layered nature of compositional challenges across multiple generalization types. The coverage principle (Type-I) explains the fundamental failure: training on “$A$ is $B$” provides no functional equivalence evidence for “$B$ is $A^{-1}$”. Property-based approaches like bidirectional training [83] partially succeed by exploiting relationship invertibility (Type-II), using architectural modifications to learn inverse mappings from the same training data. However, recent work [42] reveals deeper representational binding constraints: even though models can exploit inverse mapping properties in idealized conditions (abstract concept-level tasks), they struggle to maintain consistent concept representations when moving to realistic surface-level predictions, where entities switch roles between subject and object positions. This demonstrates how real compositional tasks involve multiple mechanism layers, where achieving one capability may be insufficient due to representational binding limitations.
Type-II generalization fundamentally differs from Type-I in that it can, in principle, transcend the limitations imposed by the coverage principle, allowing models to generalize correctly to inputs that lie outside the coverage of their training data by leveraging invariant properties of the underlying functions. Systematic characterization of pure Type-II extrapolation remains limited and is an important direction for future work.
H.3 Type-III: Shared-operator generalization
Type-III generalization emerges through reuse of identical primitive functions across different computational positions, e.g., when $f_{1}=f_{2}$ in a multi-hop reasoning task. Unlike Type-I (bounded by coverage) and Type-II (exploiting algebraic properties), Type-III generalization leverages the structural repetition of identical operations.
This mechanism is not unique to Transformers. Recurrent architectures [84] exemplify Type-III generalization through their recurrent weight sharing: the same transformation is applied at each time step, enabling the network to process sequences of varying lengths despite training on fixed-length examples. This architectural inductive bias directly encodes the assumption that the same operation should apply across temporal positions, allowing recurrent architectures to extrapolate to longer sequences [85].
In the Transformer context, Wang et al. [19] demonstrate that Universal Transformer layers [86] enable generalization beyond in-domain data in two-hop tasks, while [87] show improved compositional reasoning with weight-shared layers. These findings suggest parameter sharing as a key enabling mechanism, where a single learned implementation of a function can be applied across multiple contexts.
A prominent manifestation of Type-III generalization appears in depth extrapolation and length generalization tasks, where models generalize to deeper or longer compositional structures than seen during training [99, 100, 101, 102]. While length generalization tasks fundamentally involve applying the same computation to longer inputs (Type-III), successful generalization often requires additional mechanisms like specialized positional encodings that facilitate computation reuse across positions. The observed brittleness of current approaches [103] suggests that such inductive biases are necessary but may not be sufficient for robust systematicity.
In summary, Type-III generalization extends beyond the coverage-bound limitations of Type-I by exploiting identical operations, without requiring the algebraic property understanding characteristic of Type-II. Type-III therefore complements Types I and II, but may still operate without explicit variable binding.
H.4 Relationship to prior taxonomies of generalization
Prior taxonomies of generalization in neural networks have primarily categorized the phenomena of compositional generalization rather than the underlying mechanisms. Lake and Baroni [12] distinguish between “mix-and-match” generalization (for small training-test differences) and systematic compositional generalization (requiring abstract rule application). Other frameworks focus on empirical evaluations through benchmarks [23] or decomposing compositional abilities into constituent skills [38].
Our taxonomy differs by categorizing the mechanisms that enable generalization rather than categorizing based on the phenomena. By distinguishing between structure-based pattern matching (Type-I), algebraic property exploitation (Type-II), and shared-operator transfer (Type-III), our framework provides a lens for interpreting why models succeed or fail at certain compositional tasks. This mechanistic approach not only explains observed behaviors but also offers hypotheses about when Types II or III become necessary for certain generalization challenges.
H.5 Open questions and future directions
To operationalize this taxonomy, future work could develop diagnostics: coverage ablation tests for Type-I (systematically removing training examples to test coverage boundaries), algebraic identity probes for Type-II (testing whether models recognize properties like commutativity without explicit training), and parameter-sharing ablations for Type-III (comparing shared vs. independent weights across positions). Current examples in this paper are illustrative rather than definitive classifications; real tasks often involve combinations of mechanisms, and our taxonomy provides a scaffold for disentangling them rather than enforcing mutually exclusive categories.
Several fundamental questions remain. First, practical computations often involve mixed ingredients of generalization, spanning multiple types. When classification is ambiguous, we suggest identifying the dominant mix of types that contribute, as most realistic tasks involve combinations rather than pure instances of any single type. Second, identifying the true compositional structure underlying a task can be ambiguous, as Locatello et al. [52] demonstrate regarding multiple valid decompositions. Third, real-world datasets contain heterogeneous computational structures, and it remains unclear how neural networks accommodate this diversity. Fourth, none of the three mechanisms solves the variable binding problem, and devising diagnostics and architectures for binding remains crucial.
Future work could explore connections to complementary frameworks:
  • Minimum Description Length [73]: Type-II may compress regularities that lower description length. Similarly, Type-III can be viewed as compression through parameter reuse.
  • Formal language learning [104, 105]: Connections to grammatical inference and automata learning
  • Cognitive science [106]: Human compositional learning strategies that may inspire architectural innovations
This taxonomy is not intended as a complete classification but rather as a conceptual framework to organize our understanding of generalization mechanisms and stimulate further research into the fundamental capabilities and limitations of neural networks on compositional tasks. Ultimately, developing architectural and optimization strategies that enable neural networks to leverage all three types more robustly and ultimately addressing the variable binding problem represents an important frontier for addressing the longstanding challenge of systematicity in connectionist systems.
Appendix I Partial computation observation drives the alignment of functional equivalence representation and vocabulary space
In this section, we investigate how exposure to partial computations affects the interpretability of intermediate state representations through vocabulary space alignment. We compare two training conditions on a modified 2-Hop task with $|\mathcal{X}|=50$ and $N=10\text{k}$, after 40k epochs of training:
  1. 1.
  2. 2.
To assess interpretability, we measure the Mean Reciprocal Rank (MRR) of intermediate state representations when projected to vocabulary space using the unembedding matrix. Low MRR indicates that the model’s internal representation of intermediate state $b$ aligns with the corresponding vocabulary token.
Fig. 19 shows a striking contrast between the two conditions. Under standard training, the MRR score remains very high throughout training, indicating that intermediate representations are not aligned with vocabulary space despite the model successfully learning the compositional task. However, when partial computations are included, the MRR score becomes very high, demonstrating clear vocabulary alignment.
This finding has important implications for mechanistic interpretability research: the absence of interpretable representations through logit lens does not indicate the absence of structured internal computation. Furthermore, it suggests that interpretability techniques may need to account for how training data shapes the alignment between internal representations and vocabulary space, rather than assuming such alignment emerges naturally from task performance.

### v2

@@ G.1 Measurement Protocol for $N_{\mathrm{req}}$
To empirically determine the minimum dataset size required for reliable compositional generalization ($N_{\mathrm{req}}$), we develop a measurement protocol that accounts for practical computational constraints while ensuring robustness. For each token set size $\left|{\mathcal{X}}\right|$ and task structure, we test multiple dataset sizes until we identify the threshold point where the model successfully generalizes to the ID test set.
@@
    Measurement precision: For each identified $N_{\mathrm{req}}$, we verified that 75% of this dataset size consistently failed to meet our generalization criterion. This establishes that our measurement error is at most $\log(0.75)=0.125$ in log scale, providing confidence in the derived power-law exponents.
G.2 Measured Power-Law Scaling Exponents Across Task Structures and Model Sizes
Using our measurement protocol, we measure the required dataset size $N_{\mathrm{req}}$ across three different compositional structures (2-Hop, Parallel-2-Hop, and 3-Hop) and three model scales (68M, 96M, and 1.5B parameters). For each task structure, we vary the token set size $\left|{\mathcal{X}}\right|$ from 50 to 200, allowing us to observe the scaling relationship.
Table 2 presents the power-law exponents obtained by linear fitting $\log(\left|{\mathcal{X}}\right|)$ vs. $\log(N_{\mathrm{req}})$ plots, all with $R^{2}>0.99$. The consistency of exponents across model sizes suggests that the observed power-law scaling relates to properties of the compositional tasks themselves, rather than model capacity. This observation aligns with our theoretical derivation in Section 5.1, which predicts that the required dataset size scales at least quadratically with token set size.
Table 2: The power-law exponents for different tasks and GPT-2 sizes, obtained by linear fitting $\log(\left|{\mathcal{X}}\right|)$ vs. $\log(N_{\mathrm{req}})$ plots. $R^{2}>0.99$ for all linear fitting.
@@ G.2 Measured Power-Law Scaling Exponents Across Task Structures and Model Sizes
G.3 Robustness to Hyperparameter Variations
Figure 16: Robustness of power-law scaling relationship to hyperparameter variations in the 2-Hop task with $\left|{\mathcal{X}}\right|=50$. Each line shows the training and test accuracy curves for a different configuration: (1) baseline, (2) reduced learning rate (4e-4, half of baseline), (3) reduced weight decay (0.01, one-tenth of baseline), and (4) changed generalization criteria (test accuracy > 0.95 within 10 epochs after training accuracy > 0.95). $R^{2}>0.99$ for all linear fitting.
To verify that our observed power-law scaling relationship is not an artifact of specific hyperparameter choices, we conduct ablation studies with modified training configurations. Figure 16 demonstrates that for the 2-Hop task with $\left|{\mathcal{X}}\right|=50$, the following changes did not significantly affect the measured $N_{\mathrm{req}}$ or the derived power-law exponent:
@@ G.3 Robustness to Hyperparameter Variations
Appendix H Detailed Analysis for Non-tree Task
This section provides additional analyses that support our findings in Sec.˜7 regarding the challenges of path ambiguity in the Non-tree task.
H.1 Coverage Analysis
Figure 17: Coverage analysis for Non-tree task with $\left|{\mathcal{X}}\right|=50$. The graph shows the percentage of ID test data covered at different $k$ values across various dataset sizes (NN). Compared to the 2-Hop task (Fig.˜3, left), Non-tree has significantly lower coverage at equivalent dataset sizes, indicating that path ambiguity impedes the formation of functional equivalence relationships.
Fig.˜17 demonstrates that with equivalent training dataset sizes, a smaller percentage of ID test examples fall inside $k$-coverage for the Non-tree task compared to the 2-Hop task shown in Fig.˜3 (Left). This aligns with our theoretical analysis in Sec.˜7, which predicts that path ambiguity limits the establishment of functional equivalence relationships between input subsequences, as the model cannot generalize across different $x_{2}$ values in the Non-tree structure even when they produce the same intermediate state $b=f_{1}(x_{1},x_{2})$.
H.2 Effect of Model Scaling
Figure 18: ID test accuracy comparison between GPT-2 (96M parameters) and GPT-2-XL (1.5B parameters) on the Non-tree task with $\left|{\mathcal{X}}\right|=50$, measured 100 epochs after training accuracy exceeds 0.99. Despite the 15x increase in parameter count, the accuracy does not increase.
Fig.˜18 shows that scaling up the model size to GPT-2-XL (1.5B parameters) does not significantly improve generalization performance on the Non-tree task, even when measured 100 epochs after reaching training accuracy > 0.99. This suggests that the challenges posed by path ambiguity cannot be overcome simply by increasing model capacity, supporting our claim that the limitation is structural rather than related to model capacity.
H.3 Comparison between Mamba and GPT-2 on Non-tree Task
Figure 19: ID test accuracy comparison between GPT-2 and Mamba on the Non-tree task with $\left|{\mathcal{X}}\right|=50$, measured 100 epochs after training accuracy exceeds 0.99.
Fig.˜19 shows that the Mamba model (4 layers, hidden dimension of 256, trained with learning rate of 0.008) shows a similar trend of ID test accuracy on Non-tree task compared to GPT-2, suggesting that the generalization failure is more likely due to the task structure itself, rather than a specific model architecture.
H.4 Representation Analysis in Successful Generalization
For a model that eventually achieved near-perfect ID accuracy (0.96) after extended training (36k epochs, $\left|{\mathcal{X}}\right|=50$, $N=50\text{k}$), we conduct causal tracing analysis to understand how it achieves generalization despite path ambiguity.
@@ H.4 Representation Analysis in Successful Generalization
The causal tracing results in Fig.˜20 reveal how the model achieves generalization in the presence of path ambiguity. Across all perturbation strategies, the model’s predictions show strong causal dependence on representations at both the $x_{1}$ and $x_{2}$ positions, indicating reliance on direct access to both input tokens rather than an abstracted intermediate computation. This pattern contrasts sharply with the 2-Hop task, where causal effects concentrate primarily at positions corresponding to clustered functional equivalence representations.
@@
Appendix I Detailed Discussion on the Taxonomy for Understanding Generalization Mechanisms
In this section, we initiate a discussion to disambiguate the mixed mechanisms of generalization into isolated testable parts by sketching a preliminary taxonomy that distinguishes three complementary mechanisms of generalization. We note that we do not view our categorization as a complete one.
Type-I: Functional equivalence-based generalization (pattern matching). This is precisely what we formalized through this work: models learn that different input fragments yield identical results in shared contexts, enabling generalization to new fragment combinations. Crucially, this generalization remains bounded by coverage, and reliable generalization fails without sufficient functional equivalence evidence. In other words, it describes the ceiling of pattern matching.
Type-II: Function property-based generalization. This mechanism exploits intrinsic properties of individual primitive functions, e.g., algebraic invariances such as commutativity or input irrelevance, where certain arguments never affect the output (e.g., $f(x_{1},x_{2})=f(x_{1})$ even when distractor $x_{2}$ is present (wen2024sparse)). Unlike the previous type, this mechanism explains the generalization beyond the coverage by leveraging ‘global’ properties that hold across all possible inputs of a primitive, beyond what is actually observed. We interpret the Reversal Curse phenomenon (berglund2023reversal) as an example of the layered nature of challenges across multiple generalization types. Our framework predicts the failure of pattern matching on this problem, since training on “$A$ is BB” provides no functional equivalence evidence for “$B$ is $A^{-1}$”. An architectural modification to learn inverse mappings from the same training data to handle this problem (lv2024analysis) can be interpreted as a utilization of Type-II generalization to enable generalization beyond coverage.
Type-III: Shared-operator generalization. This mechanism emerges through the reuse of identical primitive functions across computational positions (e.g., when $f_{1}=f_{2}$). Recurrent architectures (hochreiter1997long) exemplify the utilization of this through weight sharing across time steps, enabling processing of variable-length sequences (graves2014neural). Similarly, it has been reported in Transformers with inductive biases towards reuse of the same computation through parameter sharing (dehghani2018universal; csordas2021devil; wang2024grokked) can improve generalization on complex compositional tasks where the same primitive function can be reused in various contexts. We interpret this mechanism as exploiting structural repetition.
Distinguishing mechanisms from phenomena.
Compared to prior categorizations of generalization, which focus on observed phenomena (lake2018generalization; hupkes2020compositionality), we categorize the underlying mechanisms. As noted in Sec.˜1, many behavioral studies have examined tasks mixing functional equivalence, primitives’ intrinsic properties, and operator reuse within the same benchmark, making it difficult to pinpoint the true source of success or failure. We therefore advocate clearer experimental control and community discussion around this mechanistic distinction to sharpen future analyses of neural generalization.
Implications and future directions.
Real compositional tasks typically involve combinations of all three types (and possibly more). While preliminary, we believe this taxonomy guides future research design on constructive characterization of neural networks’ generalization behaviors on discrete sequence tasks. In this broader context, this work can be understood as a characterization and formalization of pattern-matching generalization to clarify its specific boundaries. When models succeed beyond our coverage predictions, we view these as exploiting other generalization mechanisms, i.e., beyond pattern matching. Our focused study suggests that challenges to reliable generalization remain as long as models rely primarily on pattern matching, requiring methodological innovations that harness non-pattern-matching mechanisms, e.g., variable binding. We hope this preliminary taxonomy serves as a research program towards our better understanding of generalization, and confirming or refuting its utility is an empirical matter that we invite the community to explore.